OpenStudy (anonymous):

Part 1: Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a sample problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided. Part 2: View and comment on the work of at least 2 other students. Try to choose students’ whose examples are different from the ones you created.

OpenStudy (anonymous):

Upon dividing, you know that the binomial is a factor if you get a remainder of zero. It's the same principle behind the fact that \(2\) is a factor of \(4\), but not \(5\), since \(\dfrac{4}{2}=2\) and \(\dfrac{5}{2}=2+\dfrac{1}{2}\).